Calculate the molality of a solution containing 10.8 g of ethylene glycol (C2H6O2) in 360
g of water.
1
Expert's answer
2021-02-23T04:36:33-0500

n=m/Mr.

Molality= n/m

n= 10.8/62 = 0,174193548 moles.

Molality = 0,174193548/0.36 = 0,483870967 molality.